OpenStudy (anonymous):

(10x(x^3+1)-5x^2(3x^2))/(x^3+1)^2 This somehow turns into -5x^4+10x/(x^3+1)^2 Can you explain how? Thanks!

OpenStudy (anonymous):

ax(x^b) = ax^(b+1)

OpenStudy (campbell_st):

expand the brackets and then collect like terms...

OpenStudy (anonymous):

can't do it

OpenStudy (campbell_st):

ok... so we'll do it in bits... do you know the power law for multiplication..? \[x^a \times x^b = x^{a + b}\]

OpenStudy (campbell_st):

so what happens to \[10x \times x^3 + 10x \times 1 =\]

OpenStudy (anonymous):

yea I tried that on the 5x and the 3x

OpenStudy (campbell_st):

and you also need to look at \[- 5x^2 \times 3x^2 = \]

OpenStudy (campbell_st):

what would you get...?

OpenStudy (anonymous):

-15x^4

OpenStudy (campbell_st):

and the other part..?

OpenStudy (anonymous):

I dont know :/

OpenStudy (campbell_st):

\[10x^1 \times x^3 + 10x \times 1 = \]

OpenStudy (campbell_st):

can you simplify this..?

OpenStudy (anonymous):

10x^4+10x

OpenStudy (campbell_st):

ok thats good so you have  do you know about adding and subtracting like terms..? If so can you simplify the above?